OpenStudy (anonymous):

Which of the following does not factor as a perfect square trinomial? A. 16a^2-72a+81 B. 169x^2+26xy+y^2 C. x^2-18x-81 D. 4x^2+4x+1

OpenStudy (anonymous):

C

OpenStudy (anonymous):

100% sure

OpenStudy (anonymous):

\[a^2+2*a*b+b^2=\left( a+b \right)^2\] \[a^2-2*a*b+b^2=\left( a-b \right)^2 \]

OpenStudy (anonymous):

thanks